OpenStudy (anonymous):

What are the coordinates of the midpoint between (3, -7) and (-5, 2)?

OpenStudy (lgbasallote):

mid point is the middle correct? so you need to get theAVERAGE...keyword average... so to get the x coordinate of the midpoint add the two x divide by two so 3 + (-5)/2 = -2/2 = -1 y coordinate is -7 + 2 /2 = -5/2 so coordinate of midpoint is (-1, -5/2) ^_^